# WildBench: Benchmarking LLMs with Challenging Tasks from Real Users in the Wild

### v1

@@ 4 Results and Analysis
In Table 2, we present a subset of the results from our live leaderboard on Hugging Face. [^3] https://huggingface.co/spaces/allenai/WildBench For the most up-to-date results and more interactive features, such as customizing length penalties and viewing the detailed task-wise performance of each model, please refer to our live leaderboard. Our live leaderboard also supports exploring data and comparing model outputs side by side to understand the strengths and weaknesses of each model.
@@ 4.1 Leaderboard Analysis
A unique feature of the WildBench leaderboard is the ability to compare models across different task categories, which enables us to identify the strengths and weaknesses of each model on different types of tasks. In Figure 4, we select a set of popular models for analysis: Llama-3-8B-Inst [22], Llama-3-8B-Inst-SimPO [20], Yi-1.5-34B-chat [1], Llama-3-70B-Inst, GPT-4-Turbo-0409, and Claude 3 Opus. We show their performance in WB-Score across five task categories (merged from the 12 categories shown in Figure 3). Larger models like GPT-4-Turbo-0409 and Claude 3 Opus perform well across all task categories, while open LLMs like Llama-3-8B-Inst and Yi-1.5-34B-chat show weaker performance on coding and math-related tasks.
Will a 8B model outperform a 70B model?
On the AlpacaEval-2.0 leaderboard, Llama-3-8B-Inst-SimPO (LC=44.7%) significantly outperforms Llama-3-70B-Inst (LC=34.4%) [21], which is surprising and differs from our results. As shown in both Table 2 and Figure 4, our results indicate that Llama-3-8B-Inst-SimPO is generally still worse than Yi-34B-chat and Llama-3-70B-Inst. However, on information-seeking and creative tasks, Llama-3-8B-Inst-SimPO performs comparably to Llama-3-70B-Inst. Thus, we believe AlpacaEval’s evaluation results underestimate the performance of Llama-3-70B-Inst due to task selection bias in addition to the weakness of their evaluation prompting method. While the performance of Llama-3-8B-Inst-SimPO is not as good as it seems on AlpacaEval-2.0, it is indeed the best 8B model in our evaluation and outperforms some other larger models. Interestingly, Llama-3-8B-Inst-SimPO consistently improves the performance of Llama-3-8B-Inst on all task categories, resulting in a similar shape on the radar plot in Figure 4.
@@
WildBench is robust to length bias. For example, Llama-2-70B-chat and Llama-3-70B-Inst have similar output lengths (2965 vs 2983 chars), yet Llama-3-70B-Inst ranks 5th while Llama-2-70B-chat ranks 33rd on the leaderboard of 40 models. Additionally, Yi-1.5-6B’s output length is the 4th longest among the 40 models (3322 characters), but it ranks 29th on the leaderboard. This suggests that the WildBench evaluation is not biased towards longer responses, with response quality being the most important factor in the evaluation process. Additionally, we use a length penalty to ensure that longer responses are not always favored, and users can customize the length penalty to adjust the trade-off between response length and quality according to their needs. This feature is available on our live leaderboard and is illustrated in Figure 5.
4.2 Correlation to Human Judgement via ChatbotArena Elo Rating
@@ 5 Related Works
Close-ended benchmarks typically consist of multiple-choice questions and have been widely used to evaluate LLMs [3]. For example, MMLU [9] includes multi-choice questions across various subject areas. Its variants include CMMLU [13] for Chinese, KMMLU [32] for Korean, and MMLU-Pro [36] for more challenging evaluation. GPQA [31] is another close-ended benchmark designed to be challenging even for humans with internet access. Specialized benchmarks with ground-truth answers, such as GSM8K [5] and MATH [10], also fall into this category. While these benchmarks focus on close-form answers, our work evaluates LLMs’ ability to generate free-form responses and engage in conversations with users.
@@
Several open-ended generation benchmarks rely on data curated by human experts or crowdsourcing workers. For instance, MT-Bench [38] manually creates examples for predefined categories. AlpacaEval [15] is based on author-written examples [7, 33, 35], which primarily consists of simple instructions such as rewriting tasks.
@@
A key feature of our work is that its underlying data is sourced from real-world use cases, ensuring alignment with actual LLM use cases. Notable benchmarks using real-world data include ChatbotArena [38, 4], where users input their questions and choose the better response from two LLMs. However, ChatbotArena relies on extensive human feedback. ArenaHard [14] is another work that selects user queries from ChatbotArena to construct a benchmark for automatic evaluation.
@@
Evaluating open-ended generation poses challenges due to the lack of a single valid ground truth. Human evaluation, though reliable, is expensive and time-consuming. To reduce costs and enable fast evaluation, powerful LLMs are often used as judges, as seen in benchmarks like MT-Bench, AlpacaEval, ArenaHard, and our own. Evaluation methods include single-system grading, which assigns scores to individual outputs, and pairwise comparisons, which compare outputs of two systems to compute win rates. Pairwise comparisons, while more expensive, can highlight subtle differences across systems [38]. To mitigate self-selection bias where an LLM prefers its own outputs [28], we use checklists generated from multiple LLMs, similar to InfoBench [29]. In addition, we ask LLM judges generate structured explanations that enable human verification for further calibration, inspired by Just-Eval [17].
@@
Publicly available benchmarks risk contamination from LLMs trained on such data. GPQA includes a special string to help LLM developers filter out its data [31], yet indirect leakage through cited examples remains possible. To mitigate this, we reserve a subset of WildChat that is never released publicly, similar to the SEAL benchmark [^4] https://scale.com/leaderboard, which keeps its expert-curated evaluation data private. However, WildBench provides a public validation set and details the benchmark construction process for greater transparency.
@@
While our focus is on evaluating LLM capabilities, other evaluation dimensions, such as safety [19], fairness [8], agentic planning [18, 23, 16], and hallucination detection [24, 25, 11], are equally important.
6 Limitations
Language scope.
Despite the diversity of task categories included in WildBench, the current version focuses on English. This limitation excludes evaluations of LLMs in other languages.
Data leakage concerns.
To generate our leaderboard, we must send test data through APIs for models that do not have open-source weights. This process risks data leakage, as the test data might be inadvertently incorporated into model training data in the future. To alleviate this risk, we plan to continuously update the benchmark data. Thanks to the contribution of the WildChat project, there is always new user data.
Bias in source data.
The data used in WildBench is sourced from the WildChat dataset [37], which reflects the demographic distribution of its users. Consequently, any biases present in this user base are inherited by WildBench. This includes potential biases in user demographics, interests, and interaction styles.
Multi-turn query biases.
For multi-turn queries, responses for previous turns are generated by GPT-4, as we cannot counterfactually predict what users would say next if responses were generated by the model under evaluation.
@@ 7 Conclusion and Future Directions
In this work, we introduced WildBench, a benchmark designed to evaluate LLMs using real-world user queries. By continuously updating the benchmark with new examples, WildBench strives to remain relevant and reflective of the evolving capabilities of LLMs. A unique feature of WildBench data is the nature of in-the-wild user queries with natural task distribution. To evaluate LLM performance using the collected data, we introduced a CoT-like LLM-as-judge method to improve the interpretability of evaluations and reduce ambiguity. We also incorporated a length penalty method to mitigate the length bias in LLM-as-judge evaluations. Experiments show that our primary metrics, WB-Reward and WB-Score, have very strong correlations with human judgments, surpassing existing metrics such as AlpacaEval 2 and ArenaHard.
We present extensive experiments and analyses, showcasing the performance of a wide range of 40 LLMs, including both proprietary and public ones, on the WildBench benchmark. By providing a detailed breakdown of scores across different task categories, WildBench offers insights on the strengths and weaknesses of different models. By introducing WildBench, we aim to provide a realistic, dynamic, and contamination-resilient evaluation framework that accurately reflects the capabilities of LLMs. Our leaderboard at has been visited 20K times since its launch, and we will actively maintain the project for continually evaluating new LLMs with unseen tasks over time.
Appendix A Prompt Template for Pairwise Evaluation Metric WB-Reward
Appendix B Prompt Template for Individual Evaluation Metric WB-Score
## Output Format First, please output your analysis for each model response, and then summarize your assessment to three aspects: "reason A=B", "reason A>B", and "reason B>A", and finally make your choice for the final assessment. Please provide your evaluation results in the following json format by filling in the placeholders in []: ``` {     "strengths": "[analysis for the strengths of the response]",     "weaknesses": "[analysis for the weaknesses of the response]",     "score": "[1~10]" } ```
Figure 6: Distribution of the number of turns in WildBench.
Appendix C More Information on WildBench Data
The distribution of the number of turns in WildBench can be found in Figure 6. The dataset documentation, metadata, and the public subset of WildBench can be found at https://huggingface.co/datasets/allenai/WildBench/viewer/v2. We release the data under AI2’s ImpACT license as a low-risk artifact [^5] https://allenai.org/licenses/impact-lr, and we bear all responsibility in case of rights violations. We will ensure that the dataset will be available for a long time and maintain the data by continuously updating it.
Appendix D More Information on WildBench Evaluation
Our evaluation results on the public subset of WildBench can be reproduced using evaluation scripts available at https://github.com/allenai/WildBench/tree/main. We have included generation script for each model under the folder https://github.com/allenai/WildBench/tree/main/scripts, and the scripts for evaluating generations can be found at https://github.com/allenai/WildBench/tree/main/evaluation.
Appendix E Task Categories
In Section 2.2 we mentioned that tasks are cateogrized into 12 categories to enable fine-grained analysis of LLM capabilities. The definition of these task categoriezs are as follows.
@@ Appendix F Full WildBench Leaderboard
The full WildBench leaderboard as of Jun 5, 2024 can be found in Figure 5. You can view and interact with the latest results on our leaderboard on our website at https://huggingface.co/spaces/allenai/WildBench

### v2

@@ 4 Results & Analysis
In Table 2, we present a subset of the results from our live leaderboard demo. For the most up-to-date results and more interactive features, such as customizing length penalties and viewing the detailed task-wise performance of each model, please refer to our live leaderboard. Our live leaderboard also supports exploring data and comparing model outputs side by side to understand the strengths and weaknesses of each model.
@@ 4.1 Leaderboard Analysis
A unique feature of the WildBench leaderboard is the ability to compare models across different task categories, which enables us to identify the strengths and weaknesses of each model on different types of tasks. In Figure 5, we select a set of popular models for analysis: Llama-3-8B-Inst (Meta, 2023), Llama-3-8B-Inst-SimPO (Meng et al., 2024b), Yi-1.5-34B-chat (AI et al., 2024), Llama-3-70B-Inst, GPT-4-Turbo-0409, and Claude 3 Opus. We show their performance in WB-Score across five task categories (merged from the 12 categories shown in Figure 3). Larger models like GPT-4-Turbo-0409 and Claude 3 Opus perform well across all task categories, while open LLMs like Llama-3-8B-Inst and Yi-1.5-34B-chat show weaker performance on coding and math-related tasks.
Will an 8B model outperform a 70B model?
On the AlpacaEval-2.0 leaderboard, Llama-3-8B-Inst-SimPO (LC=44.7%) significantly outperforms Llama-3-70B-Inst (LC=34.4%) (Meng et al., 2024a), which is surprising and differs from our results. As shown in both Table 2 and Figure 5, our results indicate that Llama-3-8B-Inst-SimPO is generally still worse than Yi-34B-chat and Llama-3-70B-Inst. However, on information-seeking and creative tasks, Llama-3-8B-Inst-SimPO performs comparably to Llama-3-70B-Inst. Thus, we believe AlpacaEval’s evaluation results underestimate the performance of Llama-3-70B-Inst due to task selection bias in addition to the weakness of their evaluation prompting method. While the performance of Llama-3-8B-Inst-SimPO is not as good as it seems on AlpacaEval-2.0, it is indeed the best 8B model in our evaluation and outperforms some other larger models. Interestingly, Llama-3-8B-Inst-SimPO consistently improves the performance of Llama-3-8B-Inst on all task categories, resulting in a similar shape on the radar plot in Figure 5.
@@
WildBench is robust to length bias. For example, Llama-2-70B-chat and Llama-3-70B-Inst have similar output lengths (2,965 vs 2,983 chars), yet Llama-3-70B-Inst ranks 5th while Llama-2-70B-chat ranks 33rd on the leaderboard of 40 models. Additionally, Yi-1.5-6B’s output length is the 4th longest among the 40 models (3,322 characters), but it ranks 29th on the leaderboard. This suggests that the WildBench evaluation is not biased towards longer responses, with response quality being the most important factor in the evaluation process. Additionally, we use a length penalty to ensure that longer responses are not always favored, and users can customize the length penalty to adjust the trade-off between response length and quality according to their needs. This feature is available on our live leaderboard and is illustrated in Figure 6.
4.2 Correlation to Human Judgment
@@
4.3 Ablation studies and discussions.
Checklists. In our ablation study on the impact of checklists, we compared model performance with and without checklists by removing the associated parts from the prompt templates. The results indicate that incorporating checklists improves the final correlation with human preferences. Specifically, the WB-Score without checklists achieves a Pearson correlation of 0.905 (for all models), which is lower than the 0.925 correlation achieved when using checklists.
Length penalties. We experimented with different $K$ (100, 200, 500, 1000, $\inf$) in the length penalty method. We found that $K=500$ is the best choice, as it achieves the highest correlation with human judgments. This result suggests that the length penalty method is effective in mitigating the length bias in LLM evaluations.
Do multiple LLMs as judges help? How much do multiple LLMs help? We experimented with using GPT-4, Claude 3 Opus, and Mistral-Large as LLM judges. Our experiments revealed that these LLM judges produced very similar results, thereby exerting minimal influence on the final relative ranking of LLMs. Considering to reduce the cost of evaluation and faster turnaround time, we recommend using a single LLM as a judge in practice. In the future versions, we will explore more efficient ways to use multiple LLMs as judges, for example, by using different judge LLMs for different tasks that are best suited to their strengths.
Data distribution. How do we explain that WildBench has a different distribution compared to ChatbotArena’s platform but still shows a strong correlation, even better than ArenaHard? The objective of WildBench is to evaluate LLMs on challenging tasks from real users. The ArenaElo we use for comparison is derived from the hard-English split in ChatbotArena, where human users submit tasks and vote. Thus, both WildBench and ChatbotArena aim to address the same goal. While it is practically impossible to match the exact distribution of users and tasks between the two—given that WildChat users are anonymous and ChatbotArena does not publicize its data—both are sourced from real users on the web. Consequently, this represents the best possible approach for correlating our LLM ratings with human-based ratings.
Two complementary metrics: WB-Reward & WB-Score. Both metrics use checklists and a CoT-style prompt for evaluation, utilizing the same testing data. The key differences are in their methodologies: WB-Score: Evaluates each model’s outputs individually on a scale of 1-10, with detailed explanations for each score (see Appendix); WB-Reward: Compares a model’s outputs to those of three baseline models at different performance levels for a comprehensive evaluation. Pairwise evaluations can be coarse, but using three baseline models and refined pairwise choices (e.g., much better or slightly better) mitigates this. WB-Score provides a universal score comparable across models using the same evaluation templates and checklists. Additionally, WB-Score is cheaper and faster to run (10 minutes, \$5) compared to WB-Reward, which requires 3-4 times the cost due to multiple baselines. Both metrics have their strengths and weaknesses. We use both to build our official leaderboard, allowing users to choose the most suitable metrics for their experiments.
@@ 5 Related Works
Close-ended benchmarks typically consist of multiple-choice questions and have been widely used to evaluate LLMs authors (2022). For example, MMLU (Hendrycks et al., 2020) includes multi-choice questions across various subject areas. Its variants include CMMLU (Li et al., 2023a) for Chinese, KMMLU (Son et al., 2024) for Korean, and MMLU-Pro (Wang et al., 2024) for more challenging evaluation. GPQA (Rein et al., 2023) is another close-ended benchmark designed to be challenging even for humans with internet access. Specialized benchmarks with ground-truth answers, such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), also fall into this category. While these benchmarks focus on close-form answers, our work evaluates LLMs’ ability to generate free-form responses and engage in conversations with users.
@@
Several open-ended generation benchmarks rely on data curated by human experts or crowdsourcing workers. For instance, MT-Bench (Zheng et al., 2024) manually creates examples for predefined categories. AlpacaEval (Li et al., 2023b) is based on author-written examples (Dubois et al., 2023; Taori et al., 2023; Wang et al., 2022), which primarily consists of simple instructions such as rewriting tasks.
@@
A key feature of our work is that its underlying data is sourced from real-world use cases, ensuring alignment with actual LLM use cases. Notable benchmarks using real-world data include ChatbotArena (Zheng et al., 2024; Chiang et al., 2024), where users input their questions and choose the better response from two LLMs. However, ChatbotArena relies on extensive human feedback. WildVision (Lu et al., 2024) is a similar project but designed for vision language models. ArenaHard (Li et al., 2024) is another work that selects user queries from ChatbotArena to construct a benchmark for automatic evaluation.
@@
Evaluating open-ended generation poses challenges due to the lack of a single valid ground truth. Human evaluation, though reliable, is expensive and time-consuming. To reduce costs and enable fast evaluation, powerful LLMs are often used as judges, as seen in benchmarks like MT-Bench, AlpacaEval, ArenaHard, and our own. Evaluation methods include single-system grading, which assigns scores to individual outputs, and pairwise comparisons, which compare outputs of two systems to compute win rates. Pairwise comparisons, while more expensive, can highlight subtle differences across systems (Zheng et al., 2024). To mitigate self-selection bias where an LLM prefers its own outputs (Panickssery et al., 2024), we use checklists generated from multiple LLMs, similar to InfoBench (Qin et al., 2024). In addition, we ask LLM judges generate structured explanations that enable human verification for further calibration, inspired by Just-Eval (Lin et al., 2023). There are also local evaluators that can be used to evaluate LLMs with our WildBench with open-weight LLMs, such as TIGERScore (Jiang et al., 2023) and Prometheus (Kim et al., 2024).
@@
Publicly available benchmarks risk contamination from LLMs trained on such data. GPQA includes a special string to help LLM developers filter out its data (Rein et al., 2023), yet indirect leakage through cited examples remains possible. To mitigate this, we reserve a subset of WildChat that is never released publicly, which keeps its expert-curated evaluation data private. However, WildBench provides a public validation set and details the benchmark construction process for greater transparency.
@@
While our focus is on evaluating LLM capabilities, other evaluation dimensions, such as safety (Mazeika et al., 2024; Jiang et al., 2024), fairness (Gallegos et al., 2024), logical reasoning (Lin et al., 2024), agentic planning (Liu et al., 2023; Mialon et al., 2023; Lin et al., 2022), and hallucination detection (Min et al., 2023; Mishra et al., 2024; Hong et al., 2024), are equally important.
@@ 6 Conclusion and Future Directions
In this work, we introduced WildBench, a benchmark designed to evaluate LLMs using real-world user queries. An important feature of WildBench data is the nature of in-the-wild user queries with natural task distribution. To evaluate LLM performance using the collected data, we introduced a CoT-like LLM-as-judge method to improve the interpretability of evaluations and reduce ambiguity. We also incorporated a length penalty method to mitigate the length bias in LLM-as-judge evaluations. Experiments show that our primary metrics, WB-Reward and WB-Score, have very strong correlations with human judgments, surpassing existing evaluations.
We present extensive experiments and analyses, showcasing the performance of a wide range of 40 LLMs, including both proprietary and public ones, on the WildBench benchmark. By providing a detailed breakdown of scores across different task categories, WildBench offers insights on the strengths and weaknesses of different models. By introducing WildBench, we aim to provide a realistic, dynamic, and contamination-resilient evaluation framework that accurately reflects the capabilities of LLMs. We will actively maintain the project for continually evaluating new LLMs with unseen tasks over time.
Appendix
Appendix A Task Categories
In Section 2.2 we mentioned that tasks are categorized into 12 categories to enable fine-grained analysis of LLM capabilities. The definition of these task categories are as follows.
@@ Appendix A Task Categories
Figure 8: Distribution of the number of turns in WildBench.
Please note that the following links are allenai for double-blind review, which we will update after the review process. The supplementary zip file contains the source code for the evaluation scripts, the leaderboard, and the data.
Appendix B More Information on WildBench Data
The distribution of the number of turns in WildBench can be found in Figure 8. The dataset documentation, metadata, and the public subset of WildBench can be found at https://huggingface.co/datasets/allenai/WildBench/viewer/v2. We release the data under AI2’s ImpACT license as a low-risk artifact, and we bear all responsibility in case of rights violations. We will ensure that the dataset will be available for a long time and maintain the data by continuously updating it.
Appendix C More Information on WildBench Evaluation
Our evaluation results on the public subset of WildBench can be reproduced using evaluation scripts available at https://github.com/allenai/WildBench/. We have included generation script for each model under the folder https://github.com/allenai/WildBench/tree/main/scripts, and the scripts for evaluating generations can be found at https://github.com/allenai/WildBench/tree/main/evaluation.
Appendix D Prompt Template for Pairwise Evaluation Metric WB-Reward
Appendix E Prompt Template for Individual Evaluation Metric WB-Score
## Output Format First, please output your analysis for the model response, and then summarize your assessment to two aspects: "strengths" and "weaknesses"; Finally, please write down your rating for the assessment. Please provide your evaluation results in the following json format by filling in the placeholders in []: ``` {     "strengths": "[analysis for the strengths of the response]",     "weaknesses": "[analysis for the weaknesses of the response]",     "score": "[1~10]" } ```
@@ Appendix F Full WildBench Leaderboard
The full WildBench leaderboard as of Jun 5, 2024 can be found in Figure 6; The updated leaderboard as of Sept 1, 2024 can be found in Figure 7. Note that we used a new metric named WB-Elo that is based on merging WB-Reward and WB-Score to a collection of pairwise comparisons and perform Elo rating updates on top of existing LMSYS Elo rating, thus we can have a faster and more stable leaderboard update. You can view and interact with the latest results on our leaderboard on our website at https://huggingface.co/spaces/allenai/WildBench